OpenStudy (anonymous):

whats the vertex of the function y=x^2-14+13

terenzreignz (terenzreignz):

We have a formula for getting the vertex of a quadratic equation of the form \[\Large y = \color{red}ax^2 + \color{blue}bx + \color{green}c\] And that is this point: \[\huge \left(-\frac{\color{blue}b}{2\color{red}a}\quad , \quad \frac{4\color{red}a\color{green}c - \color{blue}b^2}{4\color{red}a}\right)\]

terenzreignz (terenzreignz):

And all you have to do now is plug in the appropriate values for \(\color{red}a , \color{blue}b,\) and \(\color{green}c\)

OpenStudy (anonymous):

Alternatively, you can complete the square. You'll go from \[y=ax^2+bx+c\] to \[y=a(x-h)^2+k,\] where \((h,k)\) is the vertex and \[(h,k)=\left(-\frac{\color{blue}b}{2\color{red}a}\quad , \quad \frac{4\color{red}a\color{green}c - \color{blue}b^2}{4\color{red}a}\right)\]

OpenStudy (isaiah.feynman):